OpenStudy (anonymous):

What are the zeros of the function? -x^2 - 2x + 3

OpenStudy (aravindg):

1

OpenStudy (aravindg):

sum of coefficennts =0 so 1 is definitely a root

Parth (parthkohli):

Zeros of a function are the values of the variable(s) that bring the function to 0. \( \color{Black}{\Rightarrow 0 = -x^2 - 2x + 3}\) Here equate it with zero and then find the values of \(x\).

OpenStudy (cwrw238):

you can write this as x^2 + 2x - 3 = 0 this can be factored

OpenStudy (cwrw238):

and solved

OpenStudy (anonymous):

1 and -3 if i am not wrong :)

OpenStudy (aravindg):

yep u r right

OpenStudy (cwrw238):

yep - right

OpenStudy (anonymous):

thank you :)

OpenStudy (cwrw238):

yw

Parth (parthkohli):

And thus, it's a victory of direct answering

OpenStudy (aravindg):

:(

OpenStudy (anonymous):

u can use a formula here too x={-b+-(b*b - 4ac)^1/2}/2a if u can relate ;)

OpenStudy (aravindg):

@abhijitjain1995 i dont see any reason y u need to use that

OpenStudy (aravindg):

@ParthKohli i am against that :)

OpenStudy (cwrw238):

its easier to factor

OpenStudy (aravindg):

yep @cwrw238

Parth (parthkohli):

@AravindG yay!

OpenStudy (anonymous):

i factored it manually . Any quad eqn is of the form -> ax*x + bx +c = 0 . U can have many approaches to a single problem people.

OpenStudy (aravindg):

i should add to it that some approaches gain momentum than others in specific qn and this was a beautiful example :)

OpenStudy (cwrw238):

thats right